if S varies directly as t^2, and s=150 when t=5 find S when t=4
Answer by checkley71(8403)

150/5~2=S/4~2 OR 25S=2400 OR S=2400/50 OR S=48
S/T~2 IS THE FORMULA OF THIS RELATIONSHIP.
SO WHEN S=150 & T=5 THE EQUATION IS 150/5~2 OR 150/25=6
NOW KNOWING THIS RELATIONSHIP AND THE VALUE (S) VARIES DIRECTLY WITH THE VALUE (T)
WE USE THE SAME FORMULA TO SOLVE THE PROBLEM WHEN T=4 KNOWING THE RELATIONSHIP IS THE SAME S/4~2=6 OR S=4~2*6 OR S=8*6 OR S=48.